# DepthCrafter: Generating Consistent Long Depth Sequences for Open-world Videos

### v1

@@ Abstract
Despite significant advancements in monocular depth estimation for static images, estimating video depth in the open world remains challenging, since open-world videos are extremely diverse in content, motion, camera movement, and length. We present DepthCrafter, an innovative method for generating temporally consistent long depth sequences with intricate details for open-world videos, without requiring any supplementary information such as camera poses or optical flow. DepthCrafter achieves generalization ability to open-world videos by training a video-to-depth model from a pre-trained image-to-video diffusion model, through our meticulously designed three-stage training strategy with the compiled paired video-depth datasets. Our training approach enables the model to generate depth sequences with variable lengths at one time, up to 110 frames, and harvest both precise depth details and rich content diversity from realistic and synthetic datasets. We also propose an inference strategy that processes extremely long videos through segment-wise estimation and seamless stitching. Comprehensive evaluations on multiple datasets reveal that DepthCrafter achieves state-of-the-art performance in open-world video depth estimation under zero-shot settings. Furthermore, DepthCrafter facilitates various downstream applications, including depth-based visual effects and conditional video generation.
Figure 1: We innovate DepthCrafter, a novel video depth estimation approach, by leveraging video diffusion models. It can generate temporally consistent long depth sequences with fine-grained details for open-world videos, without the need for additional information such as camera poses or optical flow.
11footnotetext:  Joint first authors.22footnotetext:  Corresponding authors.
@@ 1 Introduction
Depth estimation from monocular images or videos, serving as the bridge linking 2D observations and the 3D world, has been a long-standing fundamental problem in computer vision. It plays a crucial role in a wide range of downstream applications, such as mixed reality, AI-generated content, autonomous driving, and robotics [72, 27, 24, 26, 40, 13, 56]. The inherent ambiguity makes it extremely challenging, as the observed information from a single view is insufficient to determine the depth of a scene uniquely.
@@
With recent advances in foundation models, we have witnessed significant progress in depth estimation from monocular images [42, 12, 67, 68, 32, 17]. However, all these methods are tailored for static images, without considering the temporal information in videos. Temporal inconsistency, or flickering, would be observed when directly applying them to videos, as shown in Fig. 1. Native video depth estimation methods [41, 76, 35, 57, 63] typically optimize a temporally consistent depth sequence in 3D space from a pre-trained image depth model, with a given or learnable calibrated camera poses. Their performance is sensitive to both the proportion of dynamic content and the quality of the camera poses. Yet, videos in the open world are diverse in content, motion, camera movement, and length, making these methods hard to perform well in practice. Moreover, the required camera poses are often non-trivial to obtain in open-world videos, particularly for long videos and videos with abundant dynamic content.
@@
In this paper, we aim to generate temporally consistent long depth sequences with high-fidelity details for open-world videos, without requiring any additional information, e.g., camera poses, optical flow, etc. Observing the strong capability of diffusion models in generating various types of videos [23, 5, 3, 7, 8, 65, 4, 64], we propose a novel approach, named DepthCrafter, to leverage the video diffusion model for video depth estimation, while maintaining the generalization ability to open-world videos. We train our DepthCrafter, a video-to-depth model, from a pre-trained image-to-video diffusion model, using our compiled paired video-depth datasets, which are in two styles, i.e. realistic and synthetic, where the realistic dataset provides rich content diversity and the synthetic dataset offers precise depth details. On the aspect of temporal context, existing video diffusion models can only produce a fixed and small number of frames at a time, e.g., 25 frames in SVD [3]. However, this is often too short for open-world video depth estimation to accurately arrange depth distributions throughout the video. Considering both the respective advantages of the two-styled datasets and the requirement of variable long temporal context, we present a three-stage training strategy to progressively train certain layers of the diffusion model on different datasets with variable lengths. By doing so, we can adapt the diffusion model to generate depth sequences with variable lengths at one time, up to 110 frames, and harvest both the precise depth details and rich content diversity. To further enable estimating depth sequences for extremely long videos in the open world, we design an inference strategy to process the video in overlapped segments and seamlessly stitch them together.
@@
We extensively evaluate our DepthCrafter on multiple datasets under zero-shot settings. Both qualitative and quantitative results demonstrate that our DepthCrafter achieves state-of-the-art performance in open-world video depth estimation, outperforming existing methods by a large margin. Besides, we demonstrate that our DepthCrafter facilitates various downstream applications, including depth-based visual effects and conditional video generation. Our contributions can be summarized below:
@@
  • We design an inference strategy to segment-wisely process videos beyond 110 frames and seamlessly stitch them together, enabling depth estimation for extremely long videos.
@@ 2 Related Work
Monocular depth estimation. Deep neural networks have dominated monocular depth estimation [14, 16, 36, 1, 38, 66, 44] for their superior performance. Nevertheless, the generalization ability to diverse open-world scenes is challenging due to the limited training data. To this end, MiDaS [48] presents an affine-invariant loss for training on mixed datasets. Depth-Anything (V2) [67, 68] followed this idea and proposed to train the model on both labeled and large-scale unlabeled images, achieving good generalization ability. Marigold [32] and GeoWizard [17] leverage the diffusion priors to realize zero-shot transfer to unseen datasets. Besides, a stream of methods focus on estimating metric depth, such as ZoeDepth [2], UniDepth [46], and Metric3D [71]. However, all these methods are tailored for static images, while our work aims to estimate temporally consistent depth sequences from open-world videos.
@@
Video depth estimation. Compared to single-image depth estimation, video depth additionally requires temporal consistency. Existing methods could be categorized into two classes: test-time optimization and feed-forward prediction. Test-time optimization methods [41, 35, 76, 10] involve an optimization procedure for each video during inference, typically requiring camera poses or optical flow. This type of method usually can produce consistent video depth, but the required camera poses may limit their applicability to open-world videos. Feed-forward prediction methods directly predict depth sequences from videos [39, 59, 57, 63, 73, 60, 69, 70], e.g., DeepV2D [57] combines camera motion estimation with depth estimation, MAMO [69] leverages memory attention, and NVDS [63] introduces a plug-and-play stabilization network. However, due to the limited training data and model capacity, these methods often fail to address the in-the-wild videos with diverse content. By leveraging video diffusion priors and our designed three-stage training strategy, our method demonstrates the ability to perform open-world video depth estimation.
Video diffusion models. Diffusion models [55, 21] have achieved high-fidelity image generation results from text descriptions benefiting from web-scale aligned image-text datasets. Consequently, these models have been extended to generating videos of various types from text or images [23, 28, 22, 7, 77, 61, 3, 5, 64, 65]. Among these methods, VDM [23] presents the first results on video generation using diffusion models, while Sora [5] has shown remarkable results in this area. SVD [3] provides the popular open-source pre-trained models for image-to-video. Trained on a well-curated video dataset, SVD can generate high-quality videos and is used as the model prior for various video-related tasks. In this paper, we leverage the video diffusion model for high-fidelity consistent video depth estimation by taking the input video as the condition. Concurrent to our work, ChronoDepth [53] also explores video depth estimation with video diffusion priors. However, ChronoDepth only supports a short temporal context, i.e. 10 frames, which is insufficient to accurately arrange depth distributions throughout the video. In contrast, our method not only supports variable-length temporal context, up to 110 frames, but also can estimate depth sequences for extremely long videos.
@@ 3 Method
Given an open-world video, $\mathbf{v}\in\mathbb{R}^{T\times H\times W\times 3}$, our goal is to estimate temporally consistent depth sequences, ${\mathbf{d}}\in\mathbb{R}^{T\times H\times W}$, with fine-grained details. Considering the diversity of open-world videos in content, motion, camera movement, and length, the challenges to achieving our goal are threefold: 1.) a comprehensive understanding of video content for generalization ability; 2.) a long and variable temporal context to arrange the entire depth distributions accurately and keep temporal consistency; and 3.) the ability to process extremely long videos. As shown in Fig. 2, we tackle these challenges by formulating the video depth estimation as a conditional diffusion generation problem to model the conditional distribution $p(\mathbf{d}\,|\,\mathbf{v})$, training a video-to-depth model from a pre-trained image-to-video diffusion model through a meticulously designed three-stage training strategy with compiled paired video-depth datasets, and crafting an inference strategy to process extremely long videos through segment-wise estimation and seamless stitching.
@@ 3.1 Preliminaries of Video Diffusion Models
Diffusion models [21, 55] learn the data distribution $p(\mathbf{x})$ by a forward diffusion process to gradually noise the data to a target distribution, e.g. the Gaussian distribution, and a reverse denoising process to iteratively recover the data from the noise by a learned denoiser. In this paper, our study is conducted based on the stale video diffusion (SVD) [3], which is a famous open-source video diffusion model. SVD adopts the EDM-framework [31] for the noise schedule and denoising process. The diffusion process is achieved by adding i.i.d. $\sigma_{t}^{2}$-variance Gaussian noise to the data $\mathbf{x}_{0}\sim p(\mathbf{x})$:
@@ 3.2 Formulation with Diffusion Models
Latent space transformation. To generate high-resolution depth sequences without sacrificing computational efficiency, we adopt the framework of Latent Diffusion Models (LDMs) [50] that perform in a low-dimensional latent space, rather than the original data space. The transformation between the latent and data spaces is achieved by a Variational Autoencoder (VAE) [33], which was originally designed for encoding and decoding video frames in SVD [3]. Fortunately, we found it can be directly used for depth sequences with only a negligible reconstruction error, which is similar to the observation in Marigold [32] for image depth estimation. As shown in Fig. 2, the latent space transformation is formulated as:
@@
Conditioning on the video. SVD is an image-to-video diffusion model that generates videos conditioned on a single image. The conditional image is fed into the U-Net in two ways, i.e., concatenating its latent to the input latent, and injecting its CLIP [47] embedding to the intermediate features via cross-attention. Yet, our DepthCrafter involves the generation of depth sequences conditioned on video frames in a frame-to-frame fashion. Therefore, we adapt the conditioning mechanism in SVD to meet our video-to-depth generation task. As shown in Fig. 2, given the encoded latent of depth sequence $\mathbf{z}^{(\mathbf{d})}$ and video frames $\mathbf{z}^{(\mathbf{v})}$ from Eq. (4), we concatenate the video latent to the input noisy depth latent frame-wisely, rather than only the first frame, to condition the denoiser for generating the depth sequence. For high-level semantic information, we embed the video frames using CLIP and then inject the embeddings in a frame-to-frame manner to the denoiser via cross-attention. Compared to the original conditioning mechanism, our adapted conditioning provides more comprehensive information from the video frames to the denoiser, which significantly improves the alignment between the generated depth sequences and the video content, as well as the temporal consistency.
3.3 Training
@@
Dataset construction. To this end, we compiled paired datasets of two styles, i.e. realistic and synthetic, where the realistic dataset is large-scale and diverse, while the synthetic dataset is miniature but fine-grained and accurate. The realistic dataset is constructed from a large number of binocular videos with a wide range of scene and motion diversity. We cut the videos according to scene changes, and apply the state-of-the-art video stereo matching method, e.g., BiDAStereo [29], to generate temporally consistent depth sequences. Finally, we obtained $\sim$200K paired video-depth sequences with the length of $50-200$ frames. The synthetic dataset is a combination of the DynamicReplica [30] and MatrixCity [37] datasets, which contains $\sim$3K fine-grained depth annotations with a length of 150 frames.
@@
Challenges of variable long temporal context. Different from image depth estimation which can determine the distribution of relative depth from a single frame, the video depth estimation requires a long temporal context to arrange the depth distributions accurately for the entire video and keep the temporal consistency. Besides, the model should support variable-length estimation as the length of open-world videos may vary significantly. However, existing open-source video diffusion models can only generate a fixed small number of frames at a time, e.g., 25 frames in SVD [3]. It is non-trivial to adapt the pre-trained model to meet this requirement, as directly fine-tuning it with long sequences is memory-consuming, for example, a modern GPU with 40GB memory can only support the training of a 25-frame sequence in SVD.
Figure 3: Inference for extremely long videos. We divide the video into overlapped segments and estimate the depth sequences for each segment with a noise initialization strategy to anchor the scale and shift of depth distributions. These estimated segments are then seamlessly stitched together with a latent interpolation strategy to form the entire depth sequence. The overlapped frames are denoted as ${o}_{i}$ and their interpolation weights are denoted as ${w}_{i}$ and $1-{w}_{i}$, respectively.
@@
Three-stage training. Considering both the two-style paired datasets and the long temporal context requirement, we design a three-stage training strategy to harvest the variety of video content, the precise depth details, as well as the support for long and variable sequences. As shown in Fig. 2, we train our DepthCrafter from the pre-trained SVD in three stages. We first train it on our large realistic dataset to adapt the model to the video-to-depth generation task. The sequence length in this stage is randomly sampled from $[1,25]$ frames, such that the model can learn to generate depth sequences with variable lengths. In the second stage, we only fine-tune the temporal layers of the model still on our large realistic dataset, with the sequence length randomly sampled from $[1,110]$ frames. The reason why we only fine-tune the temporal layers is that the temporal layers are more sensitive to the sequence length while the spatial layers are already adapted to the video-to-depth generation task in the first stage, and doing so significantly reduces memory consumption compared to fine-tuning the full model. The long temporal context in this stage enables the model to precisely arrange the entire depth distributions for long and variable sequences. In the third stage, we fine-tune the spatial layers of the model on our small synthetic dataset, with a fixed sequence length of 45 frames since the model has already learned to generate depth sequences with variable lengths in the first two stages and tuning the spatial layers would not affect the temporal context. As the depth annotations in the synthetic dataset are more accurate and fine-grained, the model can learn more precise depth details in this stage. The three-stage training strategy makes our DepthCrafter capable of generating high-quality depth sequences for open-world videos with variable lengths.

### v2

@@ Abstract
Estimating video depth in open-world scenarios is challenging due to the diversity of videos in appearance, content motion, camera movement, and length. We present DepthCrafter, an innovative method for generating temporally consistent long depth sequences with intricate details for open-world videos, without requiring any supplementary information such as camera poses or optical flow. The generalization ability to open-world videos is achieved by training the video-to-depth model from a pre-trained image-to-video diffusion model, through our meticulously designed three-stage training strategy. Our training approach enables the model to generate depth sequences with variable lengths at one time, up to 110 frames, and harvest both precise depth details and rich content diversity from realistic and synthetic datasets. We also propose an inference strategy that can process extremely long videos through segment-wise estimation and seamless stitching. Comprehensive evaluations on multiple datasets reveal that DepthCrafter achieves state-of-the-art performance in open-world video depth estimation under zero-shot settings. Furthermore, DepthCrafter facilitates various downstream applications, including depth-based visual effects and conditional video generation.
Figure 1: We innovate DepthCrafter, a novel video depth estimation approach, that can generate temporally consistent long depth sequences with fine-grained details for open-world videos, without requiring additional information such as camera poses or optical flow.
11footnotetext:  Joint first authors.   Corresponding authors.
@@ 1 Introduction
Monocular depth estimation, serving as the bridge linking 2D observations and the 3D world, has been a long-standing fundamental problem in computer vision. It plays a crucial role in a wide range of downstream applications, e.g., mixed reality, AI-generated content, autonomous driving, and robotics [73, 28, 24, 27, 40, 14, 57, 72]. The inherent ambiguity makes it challenging, as the observed information from a single view is insufficient to determine the depth of a scene uniquely.
@@
With recent advances in foundation models, we have witnessed significant progress in depth estimation from monocular images [43, 13, 67, 68, 32, 18, 47, 71]. However, all these methods are tailored for static images, without considering the temporal information in videos. Temporal inconsistency, or flickering, would be observed when directly applying them to videos, as shown in Fig. 1. Existing video depth estimation methods [41, 77, 35, 58, 63] typically try to optimize temporally consistent depth sequences from a pre-trained image depth model, with a given or learnable camera poses. Their performance is sensitive to both the proportion of dynamic content and the quality of the camera poses. Yet, videos in the open world are diverse in content, motion, camera movement, and length, making these methods hard to perform well in practice. Moreover, the required camera poses are usually non-trivial to obtain in open-world videos, particularly for long videos and videos with abundant dynamic content.
@@
In this paper, we aim to generate temporally consistent long depth sequences with high-fidelity details for diverse open-world videos, without requiring any additional information. Observing the strong capability of diffusion models in generating various types of videos [23, 6, 3, 8, 9, 65, 4, 64], we propose DepthCrafter, to leverage the video diffusion model for video depth estimation, while maintaining the generalization ability to open-world videos. To train our DepthCrafter, a video-to-depth model, from a pre-trained image-to-video diffusion model, we compile paired video-depth datasets in two styles, i.e. realistic and synthetic, since the realistic dataset provides rich content diversity and the synthetic dataset offers precise depth details. On the aspect of temporal context, most existing video diffusion models can only produce a fixed and small number of frames at one time, e.g., 25 frames in Stable Video Diffusion (SVD) [3], which, however, is usually too short for open-world video depth estimation to accurately arrange depth distributions throughout the video. To enable variable long temporal context and fuse the respective advantages of the two-styled datasets, we present a three-stage training strategy to progressively train certain layers of the diffusion model on different datasets with variable lengths. By doing so, we can adapt the video diffusion model to generate depth sequences with variable lengths at one time, up to 110 frames, and harvest both the precise depth details and rich content diversity. To further support extremely long videos, we tailor an inference strategy to process the video in overlapped segments and seamlessly stitch them together.
@@
We extensively evaluate our DepthCrafter on diverse datasets, including indoor, outdoor, static, dynamic, realistic, and synthetic videos, under zero-shot settings. Both qualitative and quantitative results demonstrate that our DepthCrafter achieves state-of-the-art performance in open-world video depth estimation, outperforming existing methods by a large margin. Besides, we demonstrate that our DepthCrafter facilitates various downstream applications, including depth-based visual effects and conditional video generation. Our contributions are summarized below:
@@
  • We design an inference strategy to segment-wisely process videos beyond 110 frames and seamlessly stitch them together, facilitating depth estimation for extremely long videos.
@@ 2 Related Work
Image depth estimation. Image depth estimation aims at predicting the depth map from a single image [15, 17, 36, 1, 38, 66, 45]. However, the generalization ability to diverse open-world scenes is hindered by the limited training data. To this end, MiDaS [49] presented the affine-invariant depth representation, enabling mixed training datasets. Depth-Anything (V2) [67, 68] followed this idea and proposed to train the model on both labeled and large-scale unlabeled images, achieving good generalization ability. Marigold [32] and follow-up works [18, 42, 21] leverage the diffusion priors to realize zero-shot transfer to unseen datasets. Besides the affine-invariant depth, another stream of methods tried to estimate the absolute metric depth [2, 47, 71, 25, 5]. All these methods are tailored for static images without considering the temporal consistency, while our work aims to generate temporally consistent long depth sequences for open-world videos.
@@
Video depth estimation. Existing video depth estimation methods could be categorized into two classes: test-time optimization and feed-forward prediction. Test-time optimization methods [41, 35, 77, 11] involve an optimization procedure for each video during inference, which typically requires camera poses or optical flow. Their results are usually consistent, but the requirement of camera poses or optical flow would limit their applicability to open-world videos. Feed-forward prediction methods directly predict depth sequences from videos [39, 60, 58, 63, 74, 61, 69, 70], e.g., DeepV2D [58] combines camera motion estimation with depth estimation, MAMO [69] leverages memory attention, and NVDS [63] introduces a plug-and-play stabilization network. However, due to the limited video depth training data, these methods often fail to address the in-the-wild videos with diverse content. By leveraging the pre-trained video diffusion model and our designed three-stage training strategy, our method demonstrates a powerful ability for open-world video depth estimation.
Video diffusion models. Diffusion models [56, 22] have shown remarkable video generation ability [23, 6, 3, 8, 9, 65, 4, 64]. Among these methods, VDM [23] presents the first results on video generation using diffusion models, Sora [6] achieves impressive performance in this area, and SVD [3] provides the popular open-source models for image-to-video generation. Trained on a well-curated video dataset, SVD can generate high-quality videos and is used as the model prior for various video-related tasks. In this paper, we leverage the video diffusion model for high-fidelity consistent video depth estimation, such that the generalization ability to open-world videos can be maintained. Concurrent to our work, ChronoDepth [54] also explores video depth estimation with video diffusion priors. However, ChronoDepth only supports a short temporal context, i.e. 10 frames, which is insufficient to accurately arrange depth distributions throughout the video. In contrast, our method not only supports variable-length temporal context at one time, up to 110 frames, but also can estimate depth sequences for extremely long videos.
@@ 3 Method
Given an open-world video, $\mathbf{v}\in\mathbb{R}^{T\times H\times W\times 3}$, our goal is to estimate temporally consistent depth sequences, ${\mathbf{d}}\in\mathbb{R}^{T\times H\times W}$, with fine-grained details. Considering the diversity of open-world videos in appearance, content motion, camera movement, and length, the challenges to achieving our goal are threefold: 1.) a comprehensive understanding of video content for generalization ability; 2.) a long and variable temporal context to arrange the entire depth distributions accurately and keep temporal consistency; and 3.) the ability to process extremely long videos. As shown in Fig. 2, we tackle these challenges by formulating the video depth estimation as a conditional diffusion generation problem $p(\mathbf{d}\,|\,\mathbf{v})$. We train a video-to-depth model from a pre-trained image-to-video diffusion model through a meticulously designed three-stage training strategy with compiled paired video-depth datasets, and tailor an inference strategy to process extremely long videos through segment-wise estimation and seamless stitching.
@@ 3.1 Preliminaries of Video Diffusion Models
Diffusion models [22, 56] learn the data distribution $p(\mathbf{x})$ by a forward diffusion process to gradually noise the data to a target distribution, e.g. the Gaussian distribution, and a reverse denoising process to iteratively recover the data from the noise by a learned denoiser. In this paper, our study is conducted based on the stale video diffusion (SVD) [3], which is a famous open-source video diffusion model. SVD adopts the EDM [31] diffusion framework. The diffusion process is achieved by adding i.i.d. $\sigma_{t}^{2}$-variance Gaussian noise to the data $\mathbf{x}_{0}\sim p(\mathbf{x})$:
@@ 3.2 Formulation with Diffusion Models
Latent space transformation. To generate high-resolution depth sequences without sacrificing computational efficiency, we adopt the framework of Latent Diffusion Models (LDMs) [51] that perform in a low-dimensional latent space. The transformation between the latent and data spaces is achieved by a Variational Autoencoder (VAE) [33], which was originally designed for encoding and decoding video frames in SVD [3]. Fortunately, we found it can be directly used for depth sequences with only a negligible reconstruction error, which is similar to the observation in Marigold [32] for image depth estimation. As shown in Fig. 2, the latent space transformation is formulated as:
@@
Conditioning on the video. SVD is an image-to-video diffusion model that generates videos conditioned on a single image. The conditional image is fed into the U-Net in two ways, i.e., concatenating its latent to the input latent, and injecting its CLIP [48] embedding to the intermediate features via cross-attention. Yet, our DepthCrafter involves the generation of depth sequences conditioned on video frames in a frame-to-frame fashion. Therefore, we adapt the conditioning mechanism to meet our video-to-depth generation task. As shown in Fig. 2, given the encoded latent of depth sequence $\mathbf{z}^{(\mathbf{d})}$ and video frames $\mathbf{z}^{(\mathbf{v})}$ from Eq. (4), we concatenate the video latent to the input noisy depth latent frame-wisely, rather than only the first frame, to condition the denoiser for generating the depth sequence. For high-level semantic information, we embed the video frames using CLIP and then inject the embeddings in a frame-to-frame manner to the denoiser via cross-attention. Compared to the original conditioning mechanism, our adapted conditioning provides comprehensive information from the video frames to the denoiser, which guarantees the alignment between the generated depth sequences and the video content.
3.3 Training Strategy
@@
Dataset construction. To this end, we compiled paired datasets of two styles, i.e. realistic and synthetic, where the realistic dataset is large-scale and diverse, and the synthetic dataset is miniature but fine-grained and accurate. The realistic dataset is constructed from a large number of binocular videos with a wide range of scene and motion diversity. We cut the videos according to scene changes, and apply the state-of-the-art video stereo matching method, e.g., BiDAStereo [29], to generate temporally consistent depth sequences. Finally, we obtained $\sim$200K paired video-depth sequences with the length of $50-200$ frames. The synthetic dataset is a combination of the DynamicReplica [30] and MatrixCity [37] datasets, which contains $\sim$3K fine-grained depth annotations with a length of 150 frames.
@@
Challenges of variable long temporal context. Different from image depth estimation which can determine the distribution of relative depth from a single frame, the video depth estimation requires a long temporal context to arrange the depth distributions accurately for the entire video and keep the temporal consistency. Besides, the model should support variable-length estimation as the length of open-world videos may vary significantly. However, existing open-source video diffusion models can only generate a fixed small number of frames at a time, e.g., 25 frames in SVD [3]. It is non-trivial to adapt the pre-trained model to meet this requirement, as directly fine-tuning it with long sequences is memory-consuming, for example, modern GPUs with 40GB memory can only support the training of a 25-frame sequence in SVD.
@@
Three-stage training. Considering both the two-style paired datasets and the long temporal context requirement, we design a three-stage training strategy to harvest the variety of video content, the precise depth details, as well as the support for long and variable sequences. As shown in Fig. 2, we train our DepthCrafter from the pre-trained SVD in three stages. We first train it on our large realistic dataset to adapt the model to the video-to-depth generation task. The sequence length in this stage is randomly sampled from $[1,25]$ frames, such that the model can learn to generate depth sequences with variable lengths. In the second stage, we only fine-tune the temporal layers of the model, still on our large realistic dataset, but with the sequence length randomly sampled from $[1,110]$ frames. The reason why we only fine-tune the temporal layers is that the temporal layers are more sensitive to the sequence length while the spatial layers are already adapted to the video-to-depth generation task in the first stage, and doing so significantly reduces memory consumption compared to fine-tuning the full model. The long temporal context in this stage enables the model to precisely arrange the entire depth distributions for long and variable sequences. In the third stage, we fine-tune the spatial layers of the model on our small synthetic dataset, with a fixed sequence length of 45 frames since the model has already learned to generate depth sequences with variable lengths in the first two stages and tuning the spatial layers would not affect the temporal context. As the depth annotations in the synthetic dataset are more accurate and fine-grained, the model can learn more precise depth details in this stage. The three-stage training strategy makes our DepthCrafter capable of generating high-quality depth sequences for open-world videos with variable lengths.
Figure 3: Inference for extremely long videos. We divide the video into overlapped segments and estimate the depth sequence for each segment with a noise initialization strategy to anchor the scale and shift of depth distributions. These depth segments are then seamlessly stitched together with a latent interpolation strategy to form the entire depth sequence. The overlapped frames and their interpolation weights are denoted as ${o}_{i}$, ${w}_{i}$ and $1-{w}_{i}$, respectively.
3.4 Inference for Extremely Long Videos
Figure 4: Qualitative comparison for open-world video depth estimation. We compare with the representative single-image depth estimation method Depth-Anything-V2 [68] and the video depth estimation method NVDS [63]. For better visualizing the temporal quality, we show the temporal profiles of each result in green boxes, by slicing the depth values along the time axis at the green line positions.
4 Experiments
4.1 Implementation
We implemented our DepthCrafter based on SVD [3], using the diffusers [59] library. We train our model at the resolution of $320\times 640$ for efficiency, but we can estimate depth sequences at any resolution, e.g., $576\times 1024$, during inference. We use the Adam optimizer [34] with a learning rate of $1\times 10^{-5}$ and a batch size of 8. The number of iterations in the three stages of training is $80K$, $40K$, and $10K$, respectively. We employed eight NVIDIA A100 GPUs for training, with a total training time of about five days. Except for specific explanations, the number of denoising steps is set to 5 during inference. Please refer to the supplementary material for more implementation details.
4.2 Evaluation
Evaluation datasets. We evaluate our model on four video datasets, a single-image dataset, as well as the DAVIS dataset [46] and in-the-wild videos for qualitative results. The evaluations are conducted under the zero-shot setting, where the testing datasets cover a variety of scenes, including synthetic and realistic scenes, indoor and outdoor scenes, and static and dynamic scenes, to evaluate the generalization ability of our model across various open-world scenarios. Sintel [7] is a synthetic dataset with precise depth labels, featuring dynamic scenes with diverse content and camera motion. It contains 23 sequences with a length of around 50 frames. ScanNet v2 [12] is an indoor dataset with depth maps obtained from a Kinect sensor. We extracted 90 frames at a rate of 15 frames per second from each sequence of the test set, which includes 100 RGB-D video sequences of various scenes. Since ScanNet v2 contains only static indoor scenes, we further introduced 5 dynamic indoor RGB-D videos with a length of 110 frames each from the Bonn [44] dataset to better evaluate the performance of our model on dynamic scenes. KITTI [19] is a street-scene outdoor dataset for autonomous driving, with sparse metric depths captured by a LiDAR sensor. We adopted the validation set, which includes 13 scenes, and extracted 13 videos from it with a length of 110 frames each. Besides, we also evaluated our model for single-image depth estimation on the NYU-v2 [55] test set, which contains 654 images.
Table 1: Zero-shot depth estimation results. We compare with the best single-image depth estimation model, Marigold [32], and Depth Anything (V2) [67, 68], as well as the representative video depth estimation models, NVDS [63] and ChronoDepth [54]. Best and second best results are highlighted.